The scores on a lab test are normally distributed with mean of 200. If the standard deviation is 20, find The score that is 2 standard deviations below the mean
Answer by John10(297)

Hi,
you know the mean = 200, standard deviation = 20,
you can use the formula z = mean - (2 standard deviation)(standard deviation)
= 200 - 2(20)
= 200 - 40 = 160
John(john100185@yahoo.com)